→P Зm 2 5 Зm P 3 N, P 3m ЗР ЗР Am

determine internal force N34, N35, N36
Answer #1

Firstly we will find out reactions of supports.

Summition X =0

Therefore P+3P+P=H4 +H5

And summition y=0

Therefore V4=3P

Take moment about 5 =0,

P×3-3P×6+3P×4+H4×6-P×3=0

From heee, H4= P

Therefore, H5= 3p

Now forces in the required members are as follows,

N34= 3P(compressive)

N36=P(compressive)

For N35 we will solve the junction 5.

For that we will first find force in 1-5 member which is equal to ,

P= N15× cos(∆)

N15= P/ cos(∆)

N15= 5P/4 ( compressive)

Now take summition of forces in X direction at joint 5,

N15 × cos(∆) +N35×sin(∆)= 3P

From here N35= 10P/3 ( compressive)